A system with just two energy levels in thermal equilibrium with a heat reservior at temp. 600°k .The energy gap between the level is 0.1 eV. Find the probability that the system is in higher energy level.
1
Expert's answer
2020-03-23T13:01:40-0400
"P(1)=\frac{\exp{(-\frac{E}{kT})}}{1+\exp{(-\frac{E}{kT})}}"

"P(1)=\frac{\exp{(-\frac{0.1(1.6\cdot10^{-19})}{(1.38\cdot10^{-23})(600)})}}{1+\exp{(-\frac{0.1(1.6\cdot10^{-19})}{(1.38\cdot10^{-23})(600)})}}"

"P(1)=0.1265"